OpenStudy (anonymous):

10(6/7)-(3*6/7-11)=17 Does that equal to 17?

OpenStudy (anonymous):

Do I need to divide 6/7 ?

OpenStudy (anonymous):

it ended up being 0.8571428....

OpenStudy (anonymous):

So how do I write that ?

OpenStudy (anonymous):

\[(10 * \frac{6}{7})-(3*\frac{6}{7}-11)\] \[( \frac{60}{7})-(\frac{18}{7}-11)\] \[( \frac{60}{7})-\frac{18}{7}+11\] \[( \frac{42}{7})+11\] \[( 6+11)=17\]

OpenStudy (anonymous):

Oh !! I see it now! So you need to multiply the whole number by the fraction! Thanks so much!